OpenStudy (anonymous):

Due to Lara’s success, she is increasing the price of her signature honey lemonade from 0.20 cents a cup to 0.25 cents. What is the percent increase?

OpenStudy (anonymous):

\[\frac{0.25\, -0.2}{0.2}=\frac{0.05}{0.2}=\frac{5}{20}=\frac{1}{4}=0.25=25 \% \]